We are examining a new project. We expect to sell 6,600 units per year at $60 net cash flow apiece for the next 10 years. In other words, the annual operating cash flow is projected to be $60 × 6,600 = $396,000. The relevant discount rate is 14 percent, and the initial investment required is $1,770,000. After the first year, the project can be dismantled and sold for $1,640,000. Suppose you think it is likely that expected sales will be revised upward to 9,600 units if the first year is a success and revised downward to 5,200 units if the first year is not a success. Suppose the scale of the project can be doubled in one year in the sense that twice as many units can be produced and sold. Naturally, expansion would be desirable only if the project were a success. This implies that if the project is a success, projected sales after expansion will be 19,200. Note that abandonment is an option if the project is a failure. If success and failure are equally likely, what is the NPV of the project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) NPV $ What is the value of the option to expand?
Answer #1

A B C Initial investment $1,770,000 5 Cash Flow per unit 6 Normal units per Year 6,600 7 Number of units in case of success Number of units in case of failure Units per year after expansion 10 9 19200 -2-10 Condition Success Failure WACC Salvage Value (year1) Calculation of NPV t-0 0.5 1,770,000) $396,000 0.5 ($1,770,000)$396,000 $312,000 14% t-1 $576,000 12 13 $1,640,000 15 16 17 18 19 ithout any a ear0 10 NPV Probabili Probability NP S396,000 $576,000 $576,000 $576,000 $576,000 $1,076,587.88 0.5 $538,293.94 L20 M20 ($1,770,000) Tax wil not pass Tax will pass 0.5 $396,000$312,000 $312,000 $312,000 $312,000 ($68,887.71) 0.5($34,443.85) M23 L23 25 d NPV$503,850.08N23+N20 27 28 29 With abondonment option ear0 10 NPV Probabili Probability NP S396,000 $576,000 $576,000 $576,000 $576,000 $1,076,587.88 0.5 $538,293.94131*M31 ($1,770,000) Tax will not pass 0.5 Tax will pass 0.5 34 $2,036,000 $15,964.91 0.5 982.46 M34 L34 d NPV546,276.39N34+N31 37 39 With Expansion and abondonment option 41 ear0 10 NPV Probabili Probability NP $396,000 $1,152,000 $1,152,000 $1,152,000 $1,152,000 $3,575,807.33 0.5 $1,787,903.66 L42 M42 43 ($1,640,000) Tax will not pass 0.5 Tax will pass 0.5 $2,036,000 S15,964.91 0.5 982.46 -M45 145 47 d NPV $1,795,886.12N45+N42 50 51 52 53 Hence Expected NPV of the project is $503,850.08 Expected NPV of the project with abandonment Expected NPV of the project with $546,276.39 55expansion and abandonment $1,795,886.12 57 Value of option to exapand $1,249,609.73D55-D54Formula sheet

A B C D E F G H I J K L M N O
2
3
4 Initial investment 1770000
5 Cash Flow per unit 60
6 Normal units per Year 6600
7 Number of units in case of success 9600
8 Number of units in case of failure 5200
9 Units per year after expansion =D7*2
10
11 Condition Prob. t=0 t=1 t=2-10
12 Success 0.5 =-D4 =D6*D5 =D7*D5
13 Failure 0.5 =-D4 =D6*D5 =D8*D5
14 WACC 0.14
15 Salvage Value (year1) 1640000
16 Calculation of NPV
17 Without any abondonment option
18
19 Year0
1 2 3 10 NPV Probability Probability*NPV
20 =F12 =$G$12 =$G$12 =$G$12 =$G$12 =$C$21+G20*(1/((1+$D$14)^$G$19))+H20*PV($D$14,9,-1,0)*(1/((1+$D$14)^$G$19)) 0.5 =L20*M20 =L20*M20
21 =-D4 Tax will not pass 0.5
22 Tax will pass 0.5
23
=F12 =$G$13 =$G$13 =$G$13 =$G$13 =$C$21+G23*(1/((1+$D$14)^$G$19))+H23*PV($D$14,9,-1,0)*(1/((1+$D$14)^$G$19)) 0.5 =M23*L23 =M23*L23
24
25 Expected NPV =N23+N20 =N23+N20
26
27
28 With abondonment option
29
30 Year0
1 2 3 10 NPV Probability Probability*NPV
31 =G20 =H20 =I20 =J20 =K20 =$C$21+G31*(1/((1+$D$14)^$G$19))+H31*PV($D$14,9,-1,0)*(1/((1+$D$14)^$G$19)) 0.5 =L31*M31 =L31*M31
32 =-D4 Tax will not pass 0.5
33 Tax will pass 0.5
34
=G23+D15 =$C$21+G34*(1/((1+$D$14)^G30)) 0.5 =M34*L34 =M34*L34
35
36 Expected NPV =N34+N31 =N34+N31
37
38
39 With Expansion and abondonment option
40
41 Year0
1 2 3 10 NPV Probability Probability*NPV
42 =G20 =$D$9*$D$5 =$D$9*$D$5 =$D$9*$D$5 =$D$9*$D$5 =$C$21+G42*(1/((1+$D$14)^$G$19))+H42*PV($D$14,9,-1,0)*(1/((1+$D$14)^$G$19)) 0.5 =L42*M42 =L42*M42
43 =-D15 Tax will not pass 0.5
44 Tax will pass 0.5
45
=F13+D15 =$C$21+G45*(1/((1+$D$14)^G41)) 0.5 =M45*L45 =M45*L45
46
47 Expected NPV =N45+N42 =N45+N42
48
49
50
51 Hence
52 Expected NPV of the project is =N25
53
54 Expected NPV of the project with abandonment =N36
55 Expected NPV of the project with expansion and abandonment =N47
56
57 Value of option to exapand =D55-D54 =D55-D54
58